# MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking

### v1

@@ Abstract
Clinical decision making (CDM) is a complex, dynamic process crucial to healthcare delivery, yet it remains a significant challenge for artificial intelligence systems. While Large Language Model (LLM)-based agents have been tested on general medical knowledge using licensing exams and knowledge question-answering tasks, their performance in the CDM in real-world scenarios is limited due to the lack of comprehensive testing datasets that mirror actual medical practice. To address this gap, we present MedChain, a dataset of 12,163 clinical cases that covers five key stages of clinical workflow. MedChain distinguishes itself from existing benchmarks with three key features of real-world clinical practice: personalization, interactivity, and sequentiality. Further, to tackle real-world CDM challenges, we also propose MedChain-Agent, an AI system that integrates a feedback mechanism and a MCase-RAG module to learn from previous cases and adapt its responses. MedChain-Agent demonstrates remarkable adaptability in gathering information dynamically and handling sequential clinical tasks, significantly outperforming existing approaches. The relevant dataset and code will be released upon acceptance of this paper.
MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking
Jie Liu1thanks: J. Liu ($\spadesuit$ $\clubsuit$ $\diamondsuit$), W. Wang ($\heartsuit$ $\clubsuit$) make equal contribution. $\spadesuit$: Conceptualization; $\heartsuit$: Implementation; $\clubsuit$: Writing; $\diamondsuit$: Visualization., Wenxuan Wang2∗, Zizhan Ma2, Guolin Huang3, SU Yihang2, Kao-Jung Chang4,5, Wenting Chen1, Haoliang Li1, Linlin Shen3, Michael Lyu2 1The City University of Hong Kong 2The Chinese University of Hong Kong 3Shenzhen University 4National Yang Ming Chiao Tung University 5Taipei Veterans General Hospital Project: https://github.com/ljwztc/MedChain
@@ 1 Introduction
At the intersection of artificial intelligence and healthcare lies one of medicine’s most complex challenges: Clinical Decision Making (CDM). In healthcare delivery, CDM demands not only the synthesis of diverse data sources and continuous assessment of evolving clinical scenarios, but also evidence-based judgments for diagnosis and treatment Sutton et al. (2020). While crucial for optimal patient care, this intricate process imposes significant cognitive demands on healthcare professionals, making it an ideal candidate for AI assistance Sendak et al. (2020).
@@
Recent advances in Large Language Model (LLM)-based agents OpenAI (2023); Team et al. (2023); Gu et al. (2023); Shinn et al. (2024); Guan et al. (2023); Zhuang et al. have emerged as an effective solution for complex decision-making tasks, from software development Qian et al. (2024) to office automation Wang et al. (2024c). In the medical domain, these agents have demonstrated impressive performance on medical licensing exams Singhal et al. (2023); Pal et al. (2022) and knowledge-based assessments Gilson et al. (2023); Eriksen et al. (2023). While LLMs have consistently scored well above passing thresholds in these assessments Singhal et al. (2023), it is crucial to recognize that these assessments fall short of capturing the complexity of real-world CDM in three critical aspects.
@@
Firstly, these benchmarks rarely account for patient-specific information such as past medical history and present illness Pal et al. (2022), which significantly influence clinical decisions in real clinical scenarios. This omission fails to capture the nuanced context that often shapes personalized diagnosis. Secondly, unlike real clinical scenarios where decisions build upon previous steps, existing benchmarks present clinical tasks as independent problems Schmidgall et al. (2024), missing the critical interdependencies in the diagnostic process. In reality, clinical decision-making is a sequential process where each step is contingent upon the preceding ones, and an error in one stage can profoundly impact subsequent decisions. Thirdly, most benchmarks present all relevant information upfront, providing a static, comprehensive dataset Tu et al. (2024). However, real clinical workflow demand multiple rounds of dynamic information gathering through ongoing patient interaction.
@@
Figure 1: Demonstration of error propagation of CDM in MedChain. Starting with 2,362 initial cases, the diagram illustrates how diagnostic errors cascade through five clinical stages. Cases with incorrect diagnoses carry forward problematic information to subsequent stages, leading to a cumulative decrease in accuracy. After completing the treatment phase, we assess the correctness of decisions across all phases. Our MedChain-Agent achieves best performance in CDM comparing with other SOTA methods.
@@
MedChain: To address these critical gaps, we introduce MedChain, a novel benchmark designed to evaluate LLM-based agents in real-world clinical scenarios. Specifically, MedChain comprises 12,163 diverse cases spanning 19 medical specialties and 156 sub-categories, including 7,338 medical images with corresponding reports. Each case progresses through five crucial stages: specialty referral, history-taking, examination, diagnosis, and treatment. Unlike existing benchmarks, MedChain uniquely emphasizes three key features. 1) Personalization: Each case incorporates detailed patient-specific information. At first, agents are provided with only the patient’s chief complaint and basic information. 2) Interactivity: Information must be actively gathered through dynamic consultation from patient. 3) Sequentiality: Decisions at each stage influence subsequent steps. Only after agent independently completes all five stages, the overall diagnostic process is evaluated.
@@
  • We propose a multi-agent framework based on the characteristics of CDM. This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
  • Through extensive experiments, we demonstrate the effectiveness of MedChain and the MedChain-Agent framework in improving clinical decision-making accuracy and reliability.
MedQA 2018 1,273 Text
PubMedQA 2019 500 Text
MedMCQA 2022 193,155 Text
MMedBench 2024 53,566 Text
MedBench 2024b 300,901 Text
Asclepius 2024b 3,232 Text, Imaging
GMAI-MMBench 2024b 26,000 Text, Imaging
MedChain 12,163 Text, Imaging
@@ 2.1 Evaluation of LLM in Medcine
Benchmarking plays a vital role as a key performance indicator, directing model improvements, pinpointing weaknesses, and shaping the course of model evolution. The evaluation of LLMs in medicine has primarily focused on testing general medical knowledge through structured assessments. Leading benchmarks such as MultiMedQA Singhal et al. (2023) integrate various medical QA datasets (e.g., MedQA Jin et al. (2021), MedMCQA Pal et al. (2022)), emphasizing performance on medical licensing examination materials. Other benchmarks like PubMedQA Jin et al. (2019) focus on research-oriented queries, while several Chinese medical benchmarks Wang et al. (2023); Cai et al. (2024) evaluate models through multiple-choice questions from medical licensing exams.
While these benchmarks effectively assess general medical knowledge, they fail to capture three critical aspects of real-world clinical decision-making (see Table 1), i.e., personalization in patient care, the interactive nature of clinical consultations, and the sequential dependency of medical decisions, where each step builds upon previous findings.
@@ 2.2 LLM-based Agent in Medicine
LLM-based agents have demonstrated significant potential across various medical applications, encompassing tasks such as medical examination questions, clinical diagnoses, and treatment plans. Recent research has explored different approaches: Agent Hospital Li et al. (2024) provides medical scenario simulation, while several frameworks Yue et al. (2024); Tang et al. (2024); Kim et al. (2024b) focus on specific medical stages with multi-agent architectures. Some works target specialized aspects, such as CoD Chen et al. (2024a) for interpretable diagnostics and Ehragent Shi et al. (2024) for electronic health records (EHRs) analysis. Others, like Almanac Copilot Zakka et al. (2024), assist clinicians with EMR-specific tasks. AI Hospital Fan et al. (2024) explores interactive clinical scenarios, but it falls short in handling multi-modal medical imaging and lacks a comprehensive benchmark for evaluating multi-agent performance. To enhance these agents’ capabilities, researchers have integrated Retrieval-Augmented Generation (RAG), as demonstrated by MIRAGE Xiong et al. (2024)’s search-enhanced framework and Medical Graph RAG Wu et al. (2024)’s knowledge-based approach.
However, current approaches face two major limitations. First, existing frameworks focus on isolated medical tasks rather than providing seamless integration across different clinical stages, making them insufficient for complex scenarios requiring effective inter-stage communication Gronauer and Diepold (2022). Second, current medical RAG systems’ reliance on chunk-based indexing leads to context inconsistency and computational inefficiencies Huang and Huang (2024); Gao et al. (2024), highlighting the need for more sophisticated approaches to medical knowledge integration.
@@ 3.1 Data Collection
Quality Control. To ensure the highest standards of data integrity and clinical relevance in our benchmark, we implemented a rigorous quality control process involving a panel of five senior physicians, each with over 10 years of clinical experience. Our evaluation process examined a random sample of 6,000 cases (49.3% of the dataset). Each case was assessed on multiple dimensions including disease prevalence, clinical relevance, accuracy of patient history, appropriateness of diagnostic procedures, correctness of diagnosis, and suitability of treatment recommendations. We employed a standardized scoring system to quantify case quality across these dimensions and calculated inter-rater reliability using Cohen’s kappa coefficient. The quality assessment yielded impressive results, with an overall quality rate of 94.7% of evaluated cases meeting or exceeding our stringent quality thresholds. Dimension-specific quality rates ranged from 92.9% to 97.2%, demonstrating consistently high standards across all evaluation criteria. Our inter-rater reliability analysis produced a Cohen’s kappa coefficient of 0.82, indicating strong agreement among our expert reviewers. Cases that did not meet the thresholds (5.3%) underwent revision or were excluded from the final dataset to maintain benchmark integrity.
@@ 3.2 Clinical Workflow Simulation
The constrcution pipeline of MedChain  is sketched out here. The details of benchmark can be refer to Appendix A. We demonstrate several examples in Appendix Figure 4.
@@ 3.3 Interaction Environment
To simulate the doctor-patient consultation process in real medical scenarios, we have developed an interactive environment where the LLM-based agent under evaluation must actively gather information through dynamic interactions. We employ a local large language model, i.e., gemma2 (9b) Team et al. (2024), to serve as the patient agent. This patient agent is initialized with pre-defined case information but is unaware of the actual diagnosis, allowing it to provide symptom information and respond to inquiries in a manner that closely mimics real patient experiences. This setup ensures the interactivity for the MedChain benchmark, enabling it to evaluate the LLM’s ability to conduct effective medical consultations, ask pertinent questions, and gather crucial information for accurate diagnosis and treatment planning. The inspiration for this patient agent setup comes from standardized patients in medical education and clinical practice. Standardized patients are trained individuals who simulate real patients, allowing healthcare professionals to practice their clinical and communication skills in a controlled environment without the risks associated with real-patient interactions Barrows (1993).
@@ 3.4 Benchmark Evaluation
Given the complexity of medical decision-making, a simple binary evaluation of correct or incorrect is inadequate. To quantify the quality of CDM results, we’ve developed an evaluation system for each stage. The standard answer for each stage comprises multiple key points. The evaluated agent’s score increases with the number of standard points addressed, while irrelevant content leads to score reduction. To implement this approach, we employ the Intersection over Union (IoU) metric. For instance, in the treatment stage, the agent’s score improves as it proposes more reasonable treatment measures.
@@ 4.1 Multi-agent System with Feedback
The MedChain-Agent framework introduces a novel multi-agent system designed to simulate the complex, interconnected nature of medical decision-making processes. This system leverages the collective intelligence of multiple specialized agents, each contributing unique expertise to the overall decision-making process.
4.1.1 Agent Role
Our framework incorporates three distinct agent types, each fulfilling crucial roles in the medical decision-making pipeline:
General Agents: These specialized agents are recruited based on the specific requirements of each task. For instance, in Task 1 Specialty Referral, General Agents possess comprehensive knowledge of medical specialties and triage protocols, enabling them to effectively assess and route patients to appropriate departments. These agents engage in group discussions, simulating the collaborative nature of medical consultations.
Summarizing Agent: In each stage of the process, a summarizing agent consolidates the discussions and analyses of the general agents. This agent synthesizes the collective expertise into a coherent summary and delivers the final decision for each task, mimicking the role of a senior physician or department head within a medical context.
Feedback Agent: The feedback agent plays a crucial role in maintaining the quality and accuracy of decisions throughout the process. This agent evaluates the output of each task, providing constructive feedback and suggesting iterative improvements when necessary. By doing so, it creates a dynamic, self-correcting system that continually refines its decision-making capabilities.
4.1.2 Decision Making
With three types of agents, MedChain-Agent conducts a sophisticated decision-making process that mirrors real-world medical practice. General agents analyze patient information and engage in collaborative discussions, sharing assessments and recommendations. The summarizing agent then consolidates these insights, weighing different perspectives to formulate a preliminary decision. The feedback agent reviews this decision, evaluating its appropriateness and implications for subsequent tasks. If issues are identified, the feedback agent provides critiques and suggestions, initiating an iterative refinement process. General agents reconsider their assessments based on this feedback, while the summarizing agent refines the decision. This cycle continues until consensus is reached or a predefined number of iterations is completed, at which point the final decision is issued.

### v2

# MedChain: Bridging the Gap Between LLM Agents and Clinical Practice with Interactive Sequence

@@ Abstract
Clinical decision making (CDM) is a complex, dynamic process crucial to healthcare delivery, yet it remains a significant challenge for artificial intelligence systems. While Large Language Model (LLM)-based agents have been tested on general medical knowledge using licensing exams and knowledge question-answering tasks, their performance in the CDM in real-world scenarios is limited due to the lack of comprehensive benchmark that mirror actual medical practice. To address this gap, we present MedChain, a dataset of 12,163 clinical cases that covers five key stages of clinical workflow. MedChain distinguishes itself from existing benchmarks with three key features of real-world clinical practice: personalization, interactivity, and sequentiality. Further, to tackle real-world CDM challenges, we also propose MedChain-Agent, an AI system that integrates a feedback mechanism and a MedCase-RAG module to learn from previous cases and adapt its responses. MedChain-Agent demonstrates remarkable adaptability in gathering information dynamically and handling sequential clinical tasks, significantly outperforming existing approaches. The relevant dataset and code will be released upon acceptance of this paper.
@@ 1 Introduction
At the intersection of artificial intelligence and healthcare lies one of medicine’s most complex challenges: Clinical Decision Making (CDM). In healthcare delivery, CDM demands not only the integration of diverse data sources and continuous assessment of evolving clinical scenarios, but also evidence-based judgments for diagnosis and treatment sutton2020overview . While crucial for optimal patient care, this intricate process imposes significant cognitive demands on healthcare professionals, making it an ideal candidate for AI assistance sendak2020human .
@@
Recent advances in Large Language Model (LLM)-based agents openai2023gpt4v ; team2023gemini ; gu2023don ; shinn2024reflexion ; guan2023leveraging ; zhuangtoolchain have emerged as an effective solution for complex decision-making tasks, from software development qian2024chatdev to office automation wang2024officebench . In the medical domain, these LLMs have demonstrated impressive performance on medical licensing exams singhal2023large ; pal2022medmcqa and knowledge-based assessments gilson2023does ; eriksen2023use ; liu2025comprehensive . While LLMs have consistently scored well above passing thresholds in these benchmark singhal2023large , it is crucial to recognize that these assessments fall short of capturing the complexity of real-world CDM, where errors can cascade through multiple decision stages, as illustrated in Figure 1. Based on our analysis, CDM exhibits three key characteristics.
@@
Firstly, these benchmarks rarely account for patient-specific information such as past medical history and present illness pal2022medmcqa , which significantly influence clinical decisions in real clinical scenarios. This omission fails to capture the nuanced context that often shapes personalized diagnosis. Secondly, unlike real clinical scenarios where decisions build upon previous steps, existing benchmarks present clinical tasks as independent problems schmidgall2024agentclinic , missing the critical interdependencies in the diagnostic process. In reality, clinical decision-making is a sequential process where each step is contingent upon the preceding ones, and an error in one stage can profoundly impact subsequent decisions. Thirdly, most benchmarks present all relevant information upfront, providing a static, and comprehensive dataset tu2024towards . However, real clinical workflow demand multiple rounds of dynamic information gathering through ongoing patient interaction.
@@
Figure 1: Demonstration of error propagation of CDM in MedChain. Starting with 2,362 initial cases, the diagram illustrates how diagnostic errors cascade through five clinical stages. Cases with incorrect diagnoses carry forward problematic information to subsequent stages, leading to a cumulative decrease in accuracy. After completing the treatment phase, we count cases that maintain correctness through each consecutive phase up to the each stage. Our MedChain-Agent achieves best performance in CDM comparing with other SOTA methods.
@@
MedChain: To address these critical gaps, we introduce MedChain, a novel benchmark designed to evaluate LLM-based agents in real-world clinical scenarios. Specifically, MedChain comprises 12,163 diverse cases spanning 19 medical specialties and 156 sub-categories, including 7,338 medical images with corresponding reports. Each case progresses through five crucial stages: specialty referral, history-taking, examination, diagnosis, and treatment. Unlike existing benchmarks, MedChain uniquely emphasizes three key features. 1) Personalization: Each case incorporates detailed patient-specific information. At first, agents are provided with only the patient’s chief complaint and basic information. 2) Interactivity: Information must be actively gathered through dynamic consultation from patient. 3) Sequentiality: Decisions at each stage influence subsequent steps. Only after agent sequentially completes all five stages, the overall diagnostic process is evaluated.
@@
  • We propose a multi-agent framework based on the characteristics of CDM, called MedChain-Agent. This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
  • Through extensive experiments, we compare the performance of existing works on MedChain and the superiority of MedChain-Agent in CDM and realibility.
@@ 2.1 Evaluation of LLM in Medcine
Benchmarking plays a vital role as a key performance indicator, directing model improvements, pinpointing weaknesses, and shaping the course of model evolution. The evaluation of LLMs in medicine has primarily focused on testing general medical knowledge through structured assessments ma2025beyond . Leading benchmarks such as MultiMedQA singhal2023large integrate various medical QA datasets (e.g., MedQA jin2021disease , MedMCQA pal2022medmcqa ), emphasizing performance on medical licensing examination materials. Other benchmarks like PubMedQA jin2019pubmedqa focus on research-oriented queries, while several Chinese medical benchmarks wang2023cmb ; cai2024medbench evaluate models through multiple-choice questions from medical licensing exams. While these benchmarks effectively assess general medical knowledge, they fail to capture three critical aspects of real-world clinical decision-making (see Appendix Table 6), i.e., personalization in patient care, the interactive nature of clinical consultations, and the sequential dependency of medical decisions, where each step builds upon previous findings.
Recently, several benchmarks wu2025towards ; ouyang2024climedbench ; liu2025asclepius ; liang2024wsi ; li2025towards ; ding2025eagle have been proposed to evaluate LLMs across diverse clinical scenarios and tasks, including information extraction, text summarization, and clinical outcome prediction. However, these benchmarks primarily consist of independent question-answer pairs, where each task is evaluated in isolation. Sequential decision-making is critical in medical practice, as each patient experiences a continuous journey from initial triage through treatment to recovery. Therefore, it is essential to evaluate how LLMs perform throughout this entire clinical pipeline. Our work distinguishes itself from existing benchmarks by focusing on sequential decision-making within interactive environments, specifically evaluating LLM performance in realistic clinical scenarios that require executing the complete patient care workflow.
@@ 2.2 LLM-based Agent in Medicine
LLM-based agents have demonstrated significant potential across various medical applications wang-etal-2025-survey ; wsiagent , encompassing tasks such as medical examination questions, clinical diagnoses, and treatment plans. Recent research has explored different approaches: Agent Hospital agenthospitall provides medical scenario simulation, while several frameworks clinicalagent ; MEDAGENTS ; mdagent focus on specific medical stages with multi-agent architectures. Some works target specialized aspects, such as CoD chen2024cod for interpretable diagnostics and Ehragent shi2024ehragent for electronic health records (EHRs) analysis. Others, like Almanac Copilot zakka2024almanac , assist clinicians with EMR-specific tasks. AI Hospital fan2024ai explores interactive clinical scenarios, but it falls short in handling multi-modal medical imaging and lacks a comprehensive benchmark for evaluating multi-agent performance. To enhance these agents’ capabilities, researchers have integrated Retrieval-Augmented Generation (RAG), as demonstrated by MIRAGE MIRAGE ’s search-enhanced framework and Medical Graph RAG GraphRAG ’s knowledge-based approach.
However, current approaches face two major limitations. First, existing frameworks focus on isolated medical tasks rather than providing seamless integration across different clinical stages, making them insufficient for complex scenarios requiring effective inter-stage communication coorapation . Second, current medical RAG systems’ reliance on chunk-based indexing leads to context inconsistency and computational inefficiencies RAGsurvey1 ; RAGsurvey2 , highlighting the need for more sophisticated approaches to medical knowledge integration.
@@ 3.1 Data Collection
Quality Control. To ensure the highest standards of data integrity and clinical relevance in our benchmark, we implemented a rigorous quality control process involving a panel of five senior physicians, each with over 10 years of clinical experience. Our evaluation process examined a random sample of 6,000 cases (49.3% of the dataset). We developed a standardized scoring system that presents physicians with comprehensive case information alongside six binary quality dimensions: disease prevalence, clinical relevance, accuracy of patient history, appropriateness of diagnostic procedures, correctness of diagnosis, and suitability of treatment recommendations. Physicians evaluate each dimension through yes/no responses, with cases satisfying all dimensions considered valid.
The quality assessment yielded strong results, with 94.7% of evaluated cases meeting or exceeding our quality thresholds. Dimension-specific quality rates ranged from 92.9% to 97.2%, demonstrating consistently high standards across all evaluation criteria. Inter-rater reliability analysis produced a Cohen’s kappa coefficient of 0.82, indicating substantial agreement among our expert reviewers. Cases that failed to meet the thresholds (5.3%) were either revised or excluded from the final dataset to maintain benchmark integrity.
@@ 3.2 Clinical Workflow Simulation
The construction pipeline of MedChain  is sketched out here. The standardization process and the format of each pipeline can be refer to Appendix A. We demonstrate an example in Appendix Figure 4 and Figure 5.
@@ 3.3 Interaction Environment
To simulate authentic doctor-patient consultations, we developed an interactive environment where LLM-based agents must actively gather information through dynamic interactions. We employ Gemma2 (9b) team2024gemma as the patient agent, initializing it with pre-defined case information while withholding the actual diagnosis. Each tested LLM serves as a doctor simulator. This setup enables the agent to provide symptom information and respond to inquiries in a manner that mirrors real patient experiences, drawing inspiration from standardized patients in medical education Barrows1993AnOO . We conduct systematic evaluation with senior physician to validate the effectiveness of patient agent. The details can be found in Appendix A.2.2.
@@ 3.4 Benchmark Evaluation
Given the complexity of medical decision-making, we developed a comprehensive evaluation framework that goes beyond simple binary assessments. The standard answer for each stage comprises multiple key points. The evaluated agent’s score increases with the number of standard points addressed, while irrelevant content leads to score reduction. For each task, we employ specific evaluation metrics: Task 1 (Specialty Referral) uses accuracy and Intersection over Union (IoU) to compare predictions with ground truth. Task 2 (History-taking) calculates IoU between predicted and ground truth examination items. Task 3 leverages DocLens xie2024doclens for assessing image interpretation quality. Task 4 employs a carefully designed prompt based on National Health Commission of China guidelines to evaluate diagnostic accuracy. Task 5 measures IoU between predicted and ground truth examination items. We elaborate the evaluation metrics in Appendix A.
@@ 4.1 Multi-agent System with Feedback
The MedChain-Agent framework introduces a multi-agent system that simulates the complex, interconnected nature of medical decision-making. This system integrates specialized agents that contribute distinct expertise to the diagnostic and treatment process. We sketch the main content here and more details please refer to Appendix B.
4.1.1 Agent Roles
Our framework comprises three agent types:
General Agents: Task-specific agents recruited based on domain requirements. For specialty referral tasks, these agents possess expertise in medical specialties and triage protocols, enabling effective patient routing. They engage in collaborative discussions that mirror real-world medical consultations.
Summarizing Agent: Acts as a central coordinator, consolidating insights from general agents into coherent decisions. This agent synthesizes collective expertise and delivers final recommendations, similar to a senior physician’s role in clinical settings.
Feedback Agent: Maintains decision quality through continuous evaluation and refinement. This agent assesses outputs, provides targeted feedback, and initiates improvements when necessary, creating a self-correcting decision process.
4.1.2 Decision Making with Feedback
The decision-making process begins with general agents analyzing patient data and engaging in structured discussions. The summarizing agent then consolidates these assessments into preliminary decisions, which undergo review by the feedback agent. When issues are identified, the feedback agent initiates an iterative refinement cycle, with general agents reconsidering their assessments and the summarizing agent adjusting decisions accordingly. This process continues until reaching consensus or completing a preset number of iterations.
4.2 MedCase-RAG
To enhance decision-making capabilities, we developed MedCase-RAG, a specialized Retrieval-Augmented Generation system for medical applications. Unlike traditional medical RAG systems, our approach employs structured case representation using feature vectors that encode 12 key clinical dimensions: Age, Sex, Chief Complaint, Symptoms, Medical History, Physical Examination, Laboratory Tests, Imaging Reports, Diagnosis, Treatment Plan, Primary Department, and Secondary Department. This feature vectors is extracted by text-embedding-ada-002 from openai. When processing new cases, MedCase-RAG performs similarity searches, identifying the three most similar cases using cosine similarity metrics. This approach provides agents with comprehensive reference points for decision-making. The MedCase-RAG also features dynamic knowledge base expansion, incorporating resolved cases as pseudo-data. This mechanism enables continuous learning from new clinical experiences, adapting to evolving medical knowledge and practices.